An off-roader explores the open desert in her Hummer. First she drives 25 ∘ west of north with a speed of 7.0 km/h for 15 minutes, then due east with a speed of 13 km/h for 7.5 minutes. She completes the final leg of her trip in 22 minutes.

What is the direction and speed of travel on the final leg? (Assume her speed is constant on each leg, and that she returns to her starting point at the end of the final leg.)

What is v in km/h?
Answer #1

Speed in first leg=7 km/h

Time=7.5 minutes

Distnace travelled in first leg=7*15/60=1.75 km

Speed in second leg=13 km/h

Time in secomd leg=7.5 minutes

Distance travelled in second leg=13*7.5/60=1.625 km

Total horizontal distance=-1.75*sin25+1.625=0.885 km

Total vertical distance=1.75*cos25=1.59 km

Distance from.origin=sqrt(0.885²+1.59²)=1.82 km

Time taken for last leg=22 minutes

Speed in last leg=1.82/(22/60)=4.95 km/hr

Direction in which it will travel=arctan(0.885/1.59)=29.1° south of west